### v1

# Copyright-Aware Incentive Scheme for Generative Art Models Using Hierarchical Reinforcement Learning

Abstract.
Generative art using Diffusion models has achieved remarkable performance in image generation and text-to-image tasks. However, the increasing demand for training data in generative art raises significant concerns about copyright infringement, as models can produce images highly similar to copyrighted works. Existing solutions attempt to mitigate this by perturbing Diffusion models to reduce the likelihood of generating such images, but this often compromises model performance. Another approach focuses on economically compensating data holders for their contributions, yet it fails to address copyright loss adequately. Our approach begin with the introduction of a novel copyright metric grounded in copyright law and court precedents on infringement. We then employ the TRAK method to estimate the contribution of data holders. To accommodate the continuous data collection process, we divide the training into multiple rounds. Finally, We designed a hierarchical budget allocation method based on reinforcement learning to determine the budget for each round and the remuneration of the data holder based on the data holder’s contribution and copyright loss in each round. Extensive experiments across three datasets show that our method outperforms all eight benchmarks, demonstrating its effectiveness in optimizing budget distribution in a copyright-aware manner. To the best of our knowledge, this is the first technical work that introduces to incentive contributors and protect their copyrights by compensating them.
Copyright Protection, Incentive Scheme, Diffusion Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
Generative models are playing an increasingly important role in the web ecosystem by revolutionizing the way digital content is created and consumed (Wu et al., 2024) (Dütting et al., 2024). These models enable the automatic generation of high-quality images, audio, and video, leading to more dynamic, personalized, and engaging web experiences (Vombatkere et al., 2024). With the rise of generative models, generative art, a prominent form of artificial intelligence-generated content (AIGC), has emerged as a cutting-edge research topic (Wu et al., 2023)(Jiang et al., 2023). With the advancements in diffusion models, such as DALL·E (Ramesh et al., 2021) and Stable Diffusion (Podell et al., 2023), generative art has demonstrated remarkable progress, particularly in image generation and text-to-image tasks.
The widespread deployment of such generative art models has brought about significant challenges concerning the risk for copyright infringement. The image generative models require large amount of training data, and it’s impractical to only use non-copyrighted content to train the models. The model holder wants to leverage the high quality copyrighted data to learn the style and content and output high quality generated data. However, the generative models may generate some images that are very similar to the copyrighted training data, which may lead to copyright infringement (Ren et al., 2024).
Various methods have been developed to protect copyright in source data, with the most common strategy involving the introduction of perturbations during the training process to effectively safeguard the dataset’s copyright: (1) Unrecognizable examples (Gandikota et al., 2023; Zhang et al., 2023a) hinder models from learning essential features of protected images, either at the inference or training stages. However, this method is highly dependent on the specific images and models involved, and it lacks universal guarantees. (2) Watermarking (Dogoulis et al., 2023; Epstein et al., 2023) embeds subtle, imperceptible patterns into images to detect copyright violations, though further research is needed to improve its reliability. (3) Machine unlearning (Bourtoule et al., 2020; Huang et al., 2021; Gao et al., 2023; Nguyen et al., 2022) removes the influence of copyrighted data, supporting the right to be forgotten. (4) Dataset deduplication (Somepalli et al., 2022) helps reduce memorization of training samples, minimizing the risk of copying protected content.
Despite these efforts, existing copyright protection techniques that introduce perturbations into generative art models to safeguard data often come at the cost of output quality, reducing the model’s competitiveness. In real-world applications, generative model developers prioritize the quality of the generated content, and any compromise in this area could result in significant financial losses, potentially amounting to billions of dollars (Gong et al., 2020).
Figure 1. The Results of the Spearman’s Rank Correlation Test for Copyright Loss and Contribution indicate distinct group behaviors: Points in group A exhibit small copyright loss and significant contributions, while points in group B show minimal contribution but substantial copyright loss.
Therefore, some researchers have begun investigating economic solutions that provide fair compensation to copyright holders while maintaining the data quality of generative models. For instance, (Wang et al., 2024a) proposes treating the copyright loss of a data point as equivalent to its contribution. This approach introduces a framework where copyright owners are compensated proportionally to the contribution of their data, utilizing a royalty distribution model grounded in the probabilistic nature of AI and principles of cooperative game theory. However, In some cases, copyright loss and contribution are not the same concept. For example, some open source training data can have high contribution to the model, while the copyright loss is relatively low. On the contrary, some training data like trademarks can have low contribution to the model, while the copyright loss is relatively high when the generated content mimics some elements or idea of them. We conducted an experiment to illustrate this situation, where we use TRAK score as the contribution (see definition in Sec. 4.1) and use the proposed metric as copyright loss (see definition in Sec. 4.2). As shown in Figure 1, points in group A (red crosses) are those that have small copyright loss and big contribution to the model, while points in group B (brown triangles) are those that have big copyright loss and small contribution to the model.
Based on the above experimental observations, we hope to design an incentive scheme for the model holder, which can compensate data holders based on their contributions and copyright losses during the training process. In real-world applications, the model holder may have a total budget for training the generative model. In addition, during training, new data will be collected continuously. Due to the continuous data collection process, there would be multiple iterative versions of the generative model (Wang et al., 2024b).
In this context, we encounter several challenges when designing incentive schemes: (1) The absence of a standardized copyright metric that complies with existing copyright laws and regulations complicates the determination of whether generated images infringe on copyrights; (2) Allocating the total budget across various versions of the generated model poses difficulties. Investing excessively in earlier versions may leave new data holders, who join at a later stage, without adequate compensation.
To tackle these challenges, we introduce a Copyright-aware Incentive Scheme tailored for generative art models. Drawing inspiration from U.S. court practices (Office, 1989), which utilize a two-part test to assess copyright violations, we develop a copyright metric that aligns with these legal standards. This metric encompasses an extrinsic component that evaluates objective similarity in specific expressive elements, as well as an intrinsic component that considers subjective similarity from the perspective of a reasonable audience. By integrating both semantic and perceptual similarity, our proposed metric effectively reflects judicial criteria. Additionally, we have designed a hierarchical budget allocation method based on reinforcement learning, which first establishes the budget for each round and subsequently determines the remuneration for data holders based on their contributions and the associated copyright loss in each round.
Our contribution are as follows:
  1. (1)
    We propose the first metric for copyright loss that closely follows the procedures of copyright law.
  2. (2)
    We designed a hierarchical budget allocation method based on reinforcement learning, which first determines the budget for each round and then determines the remuneration of the data holder based on the data holder’s contribution and copyright loss in each round. This is the first incentive scheme that compensates data contributors based on copyright loss and contributions.
  3. (3)
    Extensive experimental comparisons between existing incentive schemes and our method demonstrate that our approach is more attractive to data holders who possess high-quality and original images.
2. Related works
Text-to-image Models
Generative Adversarial Networks (GAN) (Reed et al., 2016) are one of the early approaches to generating images from text. Following works like StackGAN (Zhang et al., 2017), AttnGAN (Xu et al., 2018) improve the quality and relevance of generated images of the text-to-image models. These year, diffusion models have become a prominent choice in generative modeling (Cao et al., 2024). Diffusion models are a kind of generative models that differ from the traditional models like GANs in that they iteratively transform noisy images to recover the original images through a sequence of learned denoising steps (Ho et al., 2020). (Rombach et al., 2022a) introduces Latent Diffusion Models (LDMs), which combines diffusion processes with latent variable modeling. The authors proposed a model that operates in a lower-dimensional latent space rather than pixel space. This approach significantly reduces computational costs and enables the generation of high-resolution images with less resource usage. The diffusion models have shown remarkable abilities in generating diverse and high-quality outputs.
Copyright Protection
Copyright is crucial in AI-generated images to protect intellectual property, ensure fair use, and encourage innovation by legally securing the rights of creators and developers. In both US and EU law (Office, 1989)(Parliament and Council, 2001), the substantiality of copyright infringement is one of the most important and measurable determinants. In practice, the court will perform extrinsic and intrinsic tests to measure the substantiality (Carter, 1977) and determine whether the infringement exists. There are some previous works focusing on building a generative model that avoids generating content mimic the copyrighted images. Existed method are machine unlearning, and fine-tuning. For example, ”Forgot-Me-Not” (Zhang et al., 2023b) is a method that can safely remove specified IDs, objects, or styles from a well-configured text-to-image model in as little as 30 seconds. (Vyas et al., 2023) proposes the concept ”Near Access Free(NAF)” and also a method that can output generative models with strong bounds on the probability of sampling protected content.
Reinforcement Learning
Reinforcement Learning is a popular pivotal area in machine learning, addressing the challenge of learning optimal actions through interaction with an environment. There are three basic types of RL: value-based, policy-based, and actor-critic. Q-learning is a value-based and model-free algorithm that seeks to learn the value of actions in various states, eventually deriving an optimal policy. By integrating deep learning with Q-learning, Deep Q-Network (DQN) (Mnih et al., 2015) enabled RL algorithms to handle high-dimensional state spaces, which demonstrated RL’s potential to solve complex problems previously considered intractable. In contrast to value-based methods, policy-based approaches focus directly on learning the policy itself. Recent advancements include the development of Proximal Policy Optimization (PPO) (Schulman et al., 2017), which addresses some limitations of earlier policy gradient methods. PPO improves stability and performance by using clipped objective functions and a robust optimization approach, making it highly effective for various RL tasks.
3. Preliminaries and Problem Formulation
In this section, we will introduce the preliminaries and the problem formulation. The notations used in this paper are listed in Table.1 for ease of reference.
3.1. Preliminaries
Table 1. Notations.
Symbol Description
$\mathcal{M}_{t}$ The model at round $t$
$B,B^{t}$ Total budget, the budget allocated to round $t$
$\mathcal{D}_{k}$ The dataset of data holder $k$
$\theta$ The parameter set of the model
$c^{t}_{i}$ Copyright loss of training sample $i$ at round $t$
$c^{t}_{k}$ Copyright loss of data holder $k$ at round $t$
$\beta^{t}_{i}$ Contribution of training sample $i$ at round $t$
$\beta^{t}_{k}$ Contribution of data holder $k$ at round $t$
$\hat{\beta}^{t}_{i}$ Normalized contribution of training sample $i$ at round $t$
$a,b$ Coefficients of copyright loss
$sim_{sem}$ Semantic similarity
$sim_{per}$ Perceptual similarity
$\mathcal{A}_{t}$ The action of the outer RL at round $t$
$\mathcal{R}_{t}$ The reward of the outer RL at round $t$
$r_{t}$ The reward of the inner RL at round $t$
$\lambda,\delta$ Coefficients of the reward of the inner RL
$p^{t}$ The budget allocation vector at round $t$
$P^{t}_{k}$ Payoff of the data holder $k$ at round $t$
Definition 3.1 (Text-to-image diffusion models).
Text-to-image diffusion models represent a category of algorithms that systematically introduce noise into input images to create noised samples, then reverse the process by learning to remove noise with a guided prompt embedding at each step, ultimately working towards reconstructing an image according to the prompt (Yang et al., 2024).
The forward process that transforms the input $x_{0}$ into a prior distribution, which can be written as (Podell et al., 2023):
(1) $q(\mathbf{x}_{1},\dots,\mathbf{x}_{\mathcal{T}}\mid\mathbf{x}_{0})=\prod_{t=1}^{\mathcal{T}}q(\mathbf{x}_{v}\mid\mathbf{x}_{v-1}),$
where $x_{v}$ is the sample at time step $v\in[1,\mathcal{}{\mathcal{T}}]$. The kernel $q(\mathbf{x}_{v}\mid\mathbf{x}_{v-1})$ is often defined as:
(2) $q(\mathbf{x}_{v}\mid\mathbf{x}_{v-1})=\mathcal{N}(\mathbf{x}_{v};\sqrt{1-\gamma_{v}}\mathbf{x}_{v-1},\gamma_{t}\mathbf{I}),$
where $\gamma_{v}\in(0,1)$ is a hyperparameter that controls the variance of the noise added at each time step in the diffusion process. The model can denoise the noise image $x_{v}$ and get back to $x_{0}$ by iteratively sampling from the transition kernel, where the transition kernel is defined as:
(3) $p_{\theta}(\mathbf{x}_{v-1}\mid\mathbf{x}_{v})=\mathcal{N}(\mathbf{x}_{v-1};\mu_{\theta}(\mathbf{x}_{v},v),\Sigma_{\theta}(\mathbf{x}_{v},v)),$
where $\mu_{\theta}(\mathbf{x}_{v},v)$ and $\Sigma_{\theta}(\mathbf{x}_{v},v)$ are the mean and the covariance of the Gaussian distribution, correspondingly.
We train the transition kernel so that $p_{\theta}(x_{v})$ can closely approximate the true data distribution $q(x_{0})$. Then the training objective that we want to minimize is defined as follows (Yang et al., 2024):
(4) $\mathcal{L}_{TIDM}:=\mathbb{E}_{\epsilon,v}\lVert\epsilon-\epsilon_{\theta}(z_{v},v,\mathcal{E}(l))\rVert^{2}_{2},$
where $z_{v}$ is the latent representation of the $x_{v}$ in the forward process, $\mathcal{E}$ is a text encoder for the guiding prompt input $l$.
Definition 3.2 (Data attribution method: TRAK).
TRAK (Park et al., 2023) is an efficient data attribution method to attribute both linear and non-linear models. Given a training set $D$, randomly sample $N$ subsets $\{D^{1},D^{s},...,D^{N}\}$ with a fix size $K$ from the training set $D$. The optimal model parameter set on training set $D^{s}$ is denoted as $\theta^{*}_{s}$. A random projection matrix $P$ is used in reducing the dimension.
The data attribution score for each sample $i$ using TRAK algorithm is formulated as follows (Zheng et al., 2024):
(5) $\tau_{TRAK}(x)=(\frac{1}{N}\sum_{s=1}^{N}\phi^{s}(x)^{\top}\left(\Phi^{s}{}^{\top}\Phi^{s}\right)^{-1}\Phi^{s}{}^{\top})(\frac{1}{N}\sum_{s=1}^{N}\mathcal{Q}^{s}).$
$\Phi^{s}$ is the projected gradient matrix:
(6) $\Phi^{s}=\left[\phi^{s}(x^{1});\cdots;\phi^{s}(x^{K})\right]^{\top},$
where $\phi^{s}(x)=\mathbf{P}_{s}^{\top}\nabla_{\theta}\mathcal{F}(x;\theta_{s}^{*})$ is the projected gradient.
$\mathcal{Q}^{s}$ is the weighting term:
(7) $\mathcal{Q}^{s}=\text{diag}\left(Q^{s}(x^{1}),\cdots,Q^{s}(x^{N})\right).$
where $Q^{s}(x)=\frac{\partial\mathcal{L}}{\partial\mathcal{F}}(x;\theta_{s}^{*})$ , $\mathcal{L}$ is the loss function or the training objective of the model, and $\mathcal{F}$ is the output function of the model.
3.2. Problem Formulation
Considering the generative model training process, a model holder needs data to train the generative model, a group of data holders with data $\mathcal{D}_{k}$ join the training process, and they perform the gradient updates and finetune the model.
The training time of the generative model could be long, and at each period the joining data holders can be different. The model will be updated at each round $t\in(1,T)$, and we denoted the output model after the training process of round $t$ as model version $t$ $\mathcal{M}_{t}$. Denote the training set of model $M_{t}$ by $\mathcal{D}_{t}=({\mathcal{D}}_{1},...,\mathcal{D}_{k},...,\mathcal{D}_{n})$, and the training sample $i$ by $x_{i}\in\mathcal{D}_{k}$. $\theta_{S}$ is the parameter set of the model with the training set $S$, where $S\subseteq\{\mathcal{D}_{1},...,\mathcal{D}_{k},...,\mathcal{D}_{n}\}$. The model holder has a total budget $B$, we would like to distribute the budget to the data holder $k$ in an incentive way so that the output model performance could be optimized. The problems of budget distribution naturally arise during the training process with multiple data holders: (1) how to distribute budget among the different rounds of training; (2) how to distribute budget among the data holders at each round, so that the data holders are paid fairly based on the ”contribution” and ”copyright loss” they have in the training? Our goal is to build an incentive scheme to assign a sensible payment $P^{t}_{k}$ to each data holder $k$ at round $t$ by quantifying the contribution and the copyright loss of the training sample in the training process.
To achieve our goal, we divide the whole problem into the following problems:
  1. (1)
    Contribution Assessment: The contribution of the data points is crucial in distributing the total budget $B$. The complexity of the generative models, especially diffusion models in our case, puts barriers on computing the contribution of each training data point. Traditional approaches of contribution assessment require multiple times of retraining, which is very unfeasible for big generative models like Stable Diffusion (Rombach et al., 2022b).
  2. (2)
    Copyright Loss Estimation: The copyright loss is a crucial consideration for data holders participating in model training. It is essential for the model holder to provide compensation to the data holder for this copyright loss. However, the absence of a robust metric that aligns closely with copyright law complicates the process of accurately estimating the copyright loss associated with training data
  3. (3)
    Hierarchical Budget Allocation: The total budget $B$ will be distributed in a hierarchical structure. It should first be distributed to the training of each round, then to each data holder $k$ in the model training. It’s difficult to decide an allocation that takes both the model performance and the compensation effect into account.
4. Main Approach
In this section, we firstly model the contribution and copyright loss. Then we present a hierarchical incentive scheme based on contribution and copyright loss.
4.1. Modeling Contribution
To assess the contribution of each data holder in round $t$ within the diffusion model, we can simplify the problem to analyzing the contribution of a single training sample $i$. This involves determining how the distribution of generated output images is affected when a specific training sample is removed.
We consider the TRAK algorithm mentioned in Definition. 3.2, and we need to decide the $\mathcal{L}$ to perform data attribution for the model. It’s natural to consider the training objective as the $\mathcal{L}$:
(8) $\mathcal{L}_{TIDM}=\mathbb{E}_{\epsilon,v}\lVert\epsilon-\epsilon_{\theta}(z_{v},v,\mathcal{E}(l))\rVert^{2}_{2}.$
However, as mentioned in follow-up algorithm D-TRAK (Zheng et al., 2024), using a design choice of $\mathcal{L}$ can lead to a 40% better data attribution performance than the default $\mathcal{L}:=\mathcal{L}_{TIDM}$, due to the reason that the default $\mathcal{L}$ is not the best at retaining the information of $\nabla_{\theta}\epsilon_{\theta}$. The revised $\mathcal{L}$ is formulated as follows:
(9) $\mathcal{L}_{D-TRAK}:=\mathbb{E}_{\epsilon,v}\lVert\epsilon_{\theta}(x_{v},v,\mathcal{E}(l))\rVert^{2}_{2}.$
In this case, the data attribution score will be formulated as:
(10) $\tau_{\text{D-TRAK}}(x)=\frac{1}{N}\sum_{s=1}^{N}\hat{\phi}^{s}(x)^{\top}\left(\left(\Phi_{\text{D-TRAK}}^{s}\right)^{\top}\Phi_{\text{D-TRAK}}^{s}\right)^{-1}\left(\Phi_{\text{D-TRAK}}^{s}\right)^{\top}$
Compared to the original TRAK formula 5, $\phi^{s}(x)=\mathbf{P}_{s}^{\top}\nabla_{\theta}\mathcal{F}(x;\theta_{s}^{*})$ is changed to $\hat{\phi}^{s}(x)=\mathbf{P}_{s}^{\top}\nabla_{\theta}\mathcal{L}_{D-TRAK}$, and correspondingly $\Phi^{s}$ is changed to $\Phi^{s}_{\text{D-TRAK}}=\left[\hat{\phi}^{s}(x^{1});\cdots;\hat{\phi}^{s}(x^{K})\right]^{\top}$; $\mathcal{Q}^{s}$ is reduced to the identity matrix $I$ since we assume $\mathcal{L}$ and $\mathcal{F}$ are the same (Zheng et al., 2024).
We use the above data attribution method in the text-to-image diffusion model and compute the score of each training sample as their contribution. The contribution of sample $i$ is formulated as:
(11) $\beta_{i}=\tau_{D-TRAK}(x_{i})$
We denote the contribution of training sample $i$ as $\beta^{t}_{i}$, the minimum and maximum of the contribution in the current round $t$ as $\beta^{t}_{min}$ and $\beta^{t}_{max}$, respectively, then, we normalize the raw contribution of each training sample as:
(12) $\hat{\beta^{t}_{i}}=\frac{\beta^{t}_{i}-\beta^{t}_{min}}{\beta^{t}_{max}-\beta^{t}_{min}}$
The contribution of each data holder $k$ is computed as:
(13) $\beta^{t}_{k}=\sum_{i}\hat{\beta^{t}_{i}}$
4.2. Modeling Copyright Loss
Figure 2. Workflow of Computation of Copyright Loss.
Estimating copyright loss presents a significant challenge due to the difficulty in quantifying copyright infringement. While some previous studies have addressed copyright-related issues, they often lack explicit definitions of copyright infringement, instead using vague terms such as replications (Somepalli et al., 2022) and memorization (Carlini et al., 2023). For instance, (Vyas et al., 2023) introduces a stable algorithm as a surrogate for addressing copyright concerns, but it falls short in tracking copyright loss at the instance level for each training sample. Our goal is to propose a metric that accurately quantifies copyright infringement and assesses copyright loss associated with each individual sample.
According to US law, courts consider four factors when assessing fair use defenses (Office, 1989): (1) the purpose of the challenged use, including whether the use is of a commercial nature or is for nonprofit educational purposes, (2) the nature of the copyrighted works, (3) the amount and substantiality of the taking, and (4) the effect of the challenged use on the market for or value of the copyrighted work. From technical aspect, the substantiality will be the most measurable and relative determinants. The induced question will be: how to measure substantiality? In practice, the court will perform extrinsic and intrinsic tests to measure the substantiality (Carter, 1977). The extrinsic test is an objective comparison of the expressive elements in the works. Then, intuitively, we need a metric that measures the similarity of the elements between original images and generated images. The intrinsic test is generally conducted by the jury, and it’s related to the vision perception of human being and subjective feelings. This inspires us to propose metrics to measure substantiality from two aspects: (1) semantic metric and (2) perceptual metric.
Figure 3. Workflow of Hierarchical Reinforcement Learning. The outer RL distributes the budget $B^{1},B^{2},...,B^{T}$ to each round. The inner RL distributes the budget to each data holder according to the contribution and the copyright loss, and the model with be trained with the corresponding data. The quality of the output model serves as the reward of the outer RL.
Semantic Similarity
The semantic metric measures the similarity of the extrinsic elements between the original image and the generated image. It is inspired by the extrinsic test in the court, and it tries to capture the similarity between the original image and the generated one. Following (Ma et al., 2024), we use CLIP (Radford et al., 2021) to generate the semantic embedding for each training sample and the generated sample, and compute the MSE for measurement. Denote the generated semantic embedding of training sample $s_{i}$ and the generated image by $CLIP(s_{i})$ and $CLIP(y_{i})$, respectively, then,
(14) $sim_{sem}=MSE(CLIP(s_{i})-CLIP(y_{i}))$
Perceptual Similarity
We choose the perceptual metric as one of the measurement because it tries to mimic the vision perception of human being, and it aligns with the ”the subjective comparison” by ”reasonable audience” in the intrinsic test of the courts (Carter, 1977). Following the framework of (Zhang et al., 2018), we can compute the perceptual similarity of the training sample and the generated image. We compute deep embeddings with a network $\mathbf{F}$, then normalize the activations in the dimension of channel. Next, we scale the features with a weight and compute the average across the $l_{2}$ distance in the spatial dimensions, considering all layers.
(15) $sim_{per}=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}||w_{l}\odot(\hat{z_{i}}-\hat{y_{i}})||_{2}^{2},$
where $\hat{z_{i}}$ and $\hat{y_{i}}$ are the deep embeddings of the training sample and the generated sample got from a deep network, $w_{l}$ is the scaled vector at layer $l$, $H_{l}$ and $W_{l}$ are the height and width of the feature map at layer $l$.
Then, the copyright loss of sample $i$ is defined as:
(16) $c_{i}=asim_{sem}+bsim_{per}$
where $a$ and $b$ are hyperparameters that control the importance attached to each measurement.
Workflow Now we can compute the copyright loss with the metrics defined above for each training sample with respect to each generated image. As shown in Figure. 2, the workflow for this part is as follows: initially, each training sample is passed through the CLIP model (Radford et al., 2021) to generate captions, which are fed into a latent diffusion model to generate corresponding images. Next, we calculate the semantic similarity $sim_{sem}$ using Equation 14. Following this, in line with the framework proposed by Zhang et al. (Zhang et al., 2018), we compute the perceptual similarity $sim_{per}$. Once we have obtained both $sim_{sem}$ and $sim_{per}$, we proceed to calculate the raw copyright score $c^{t}_{i}$ for each sample $i$ at iteration round $t$. Finally, the total copyright loss for data holder $k$ at round $t$ is determined by summing the individual scores across all samples, as expressed in the equation below:
(17) $c^{t}_{k}=\sum_{i}c^{t}_{i}$
Algorithm 1 Hierarchical Budget Allocation
1:Input: Total budget $B$; pre-trained model $\mathcal{M}$; the data holders $\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{n}$; round number $T$.
2:Output: Trained model $\mathcal{M}_{T}$, the budget allocation at each round $B^{1},B^{2},...,B^{T}$.
3:Outer Budget Allocation:
4: Initialize Q-network with random weights
5: Initialize replay buffer $\mathcal{D}$
6: for each time slot $t=1$ to $T$ do
7:     Observe state $\mathcal{S}_{t}=(N_{t},C_{t},X_{t},B_{lt},t_{l})$
8:     Select action $\mathcal{A}_{t}=B^{T}$ based on $\epsilon$-greedy policy from Q-network
9:     Inner Budget Allocation:
10:     Initialize Q-network for inner allocation with random weights
11:     Initialize replay buffer $\mathcal{D}_{inner}$
12:     if $t=1$ then
13:          Load the pre-trained model $\mathcal{M}$
14:    end if
15:     for iteration $j=1$ to $10^{5}$ do
16:          Compute the contribution and copyright of each images $i$
17:          Select action $\mathbf{p}^{t}=(p^{t}_{1},p^{t}_{2},\ldots,p^{t}_{n})$, where each $p^{t}_{k}$ represents the fraction of the budget distributed to data holder $k$
18:          Distribute budget: $P^{t}=B^{t}\times\mathbf{p}^{t}$
19:          Compute rewards: $r_{t}=\lambda R_{con}-\delta R_{cop}$, where: $R_{con}=\sum_{k}{p_{k}}\times\hat{\beta_{k}},\quad R_{cop}=\sum_{k}{p_{k}}\times c_{k}$
20:          Store transition $(B^{t},\mathbf{p}^{t},r_{t})$ in $\mathcal{D}_{inner}$
21:          Sample mini-batch from $\mathcal{D}_{inner}$
22:          Update Q-network for inner allocation
23:    end for
24:     Train the model with selected data
25:     Receive outer RL reward $\mathcal{R}_{t}$
26:     if $t_{l}>1$ then
27:          $\mathcal{R}_{t}=0.01$
28:     else
29:          $\mathcal{R}_{t}=Q(\mathcal{M})$triangleright Model quality measured by FID score
30:    end if
31:     Store transition $(\mathcal{S}_{t},\mathcal{A}_{t},\mathcal{R}_{t},\mathcal{S}_{t+1})$ in $\mathcal{D}$
32:     Sample mini-batch from $\mathcal{D}$
33:     Update Q-network
34:end for
4.3. Hierarchical Budget Allocation
The budget allocation problem could be divided into distributing the budget across time slots and within a specific time slot among data holders. To deal with this issue, we use the hierarchical RL as the backbone framework as it leverages the adaptive nature of RL and is able to handle the complexity in different stages (Hutsebaut-Buysse et al., 2022).
4.3.1. Outer Budget Allocation
The outer budget allocation in the hierarchical budget allocation method is used to allocation the total budget $B$ among time slots $t$, i..e, training rounds of the model. It is developed based on the DQN framework (Li, 2017) and detailed as:
State: The state of the outer budget allocation is composed of 1) historical information of the previous rounds, including the amount of data points contributed by participating data holders $N_{t}$, the total copyright loss in the previous training up to time $t$, denoted as $C_{t}$, the total contribution in the previous training up to time $t$, denoted as $X_{t}$ 2) the leftover budget till time slot $t$, denoted as $B_{lt}$, and 3) the maximum leftover time slot $t_{l}$:
(18) $\mathcal{S}_{t}=(N_{t},C_{t},X_{t},B_{l,t},t_{l})$
Action: The action of the outer budget allocation is the budget for the current round, denoted as $\mathcal{A}_{t}$.
Reward: The reward $\mathcal{R}_{t}$ is formulated as the quality of the generative model measured by the Fréchet Inception Distance (FID) score (Heusel et al., 2018):
(19) $\mathcal{R}_{t}=\left\{\begin{array}[]{rcl}0.01&&{ift_{l}>1}\\ Q(\mathcal{M})&&{otherwise}\end{array}\right.$
The $Q(\mathcal{M})$ is the output quality of the model (See Equation. 23). The design of the reward $R_{t}$ means that at the last round $T$, the reward will be the model quality measured by the metric 23; at the middle rounds $t<T$, we will attribute a small amount of reward to encourage the budget distribution process.
4.3.2. Inner Budget Allocation
The inner budget allocation aims to allocate the round budget $B^{T}$, i.e., $\mathcal{A}_{t}$ described above, among the various data holders participating in the corresponding training round. Specifically, similar to the outer budget allocation, the inner budget allocation is also grounded in a DQN framework:
State: The state of the inner budget allocation is round budget $B^{T}$, i.e., the action of the outer budget allocation.
Action: The action of the inner budget allocation is denoted as a vector $p^{t}$, recording the fraction of the budget distributed to different data holders at the time slot $t$.
(20) $\mathbf{p^{t}}=\begin{pmatrix}p^{t}_{1}&p^{t}_{2}&\cdots&p^{t}_{k}&\cdots&p^{t}_{n}\end{pmatrix},$
where $p_{1}+p_{2}+...+p_{n}=1$. Then, the payment for each data holder $k$ in round $t$, denoted as $P^{t}_{k}$ can be computed by the fraction:
(21) $P^{t}_{k}=B^{t}\times p^{t}_{k}$
Reward: The reward is formulated as:
(22) $r_{t}=\lambda R_{con}-\delta R_{cop},$
where $R_{con}=\sum_{k}{p_{k}}\times\hat{\beta_{k}}$ and $R_{cop}=\sum_{k}{p_{k}}\times c_{k}$. The reward is based on the intuition that data holder with high contribution and high copyright should be given high payoffs. We design the reward to help the system pick the data holders with the larger contribution and the lower the copyright loss. The setting can help the system spend the budget in a smart way that allocate more payoff to the data holders that make greater contribution to the training while doesn’t need to compensate a lot for the copyright loss.
4.3.3. Workflow
Figure 3 shows the workflow of the Hierarchical Budget Allocation proposed in our method, which is described as:
  1. (1)
    At round $t$, the agent of the outer layer RL allocates the budget $B^{t}$ to the inner layer.
  2. (2)
    Train the model with data in all data holders. Compute the contribution of each training sample $\hat{\beta}^{t}_{i}$ and the copyright loss of each training sample $c^{t}_{i}$.
  3. (3)
    The agent of the inner RL get the reward $r$ of the inner layer RL according to Equation. (22).
  4. (4)
    The agent of the inner layer RL allocates the budget $p^{t}_{i}$ to each data holders with a decided portion $p_{t}$.
  5. (5)
    The agent of the inner RL get a new reward $r$ of the inner layer RL according to Equation. (22), and repeat the process until the stopping step.
  6. (6)
    The inner layer RL finishes training and output the final selection of the data holders joining the training.
  7. (7)
    The model is trained with the selected data holders at the round t.
  8. (8)
    Repeat until the final round is reached and compute the model quality of the final model according to Equation. (23). The value will be the reward of the outer layer RL.
  9. (9)
    Repeat the above process until the stopping step of the outer RL is reached. Output the final model $M_{T}$ and the budget allocation.
The proposed Hierarchical Budget Allocation method is summarized in Algorithm. 1.
Table 2. Model Output Quality for Comparative Experiments on Three Datasets
Dataset Ours G+L L+L R+R RL+R RL+L R+RL L+RL G+RL
ArtBench 6.178$\pm$0.344 5.065$\pm$0.161 3.568$\pm$0.020 3.842$\pm$0.366 5.692$\pm$0.470 5.782$\pm$0.114 5.017$\pm$0.609 4.259$\pm$0.122 5.993$\pm$0.279
Portrait 8.331$\pm$1.785 5.767$\pm$0.106 5.986$\pm$0.081 6.017$\pm$0.966 8.169$\pm$2.011 8.186$\pm$1.257 6.057$\pm$1.397 7.509$\pm$0.425 5.943$\pm$0.195
Cartoon 6.038$\pm$0.737 5.032$\pm$0.644 3.955$\pm$0.211 4.508$\pm$1.307 5.613$\pm$0.707 5.947$\pm$0.520 4.594$\pm$0.973 4.219$\pm$0.413 4.516$\pm$0.529
(a) ArtBench
(b) Portrait
(c) Cartoon
Figure 4. Comparison of Our Method with the Baselines on Three Datasets.
5. Experiments
5.1. Experiment Setup
Dataset. We use the following three datasets:
  1. (1)
    ArtBench: ArtBench(Liao et al., 2022) includes 60000 selected high-quality artworks in 10 distinctive art styles with 256x256 resolution. We gathered 800 256x256 images from the Impressionism subset.
  2. (2)
    Portrait: The Portrait dataset(Ma et al., 2024) includes 200 images of celebrities from Wikipedia with detailed annotations about the outfit, facial features and gesture of the people. The copyright of a portrait encompasses an individual’s authority over their own image, including their facial features, likeness, and posture.
  3. (3)
    Cartoon: The cartoon dataset(Ma et al., 2024) includes 200 images of cartoon characters from animations and cartoons with detailed notations. The images are selected from Wikipedia [^1] https://www.tensorflow.org/datasets/catalog/wikipedia. These cartoon images, featuring unique graphic expressions, are often considered copyrighted due to their distinctive characteristics.
Environments and models. We conducted the experiments on 4 NVIDIA A-100. The diffusion model we used is one of the newest models in Stable Diffusion series: Stable Diffusion XL 1.0222https://github.com/huggingface/diffusers. The inner and outer layer of hierarchical RL are modified from the DQN framework based on the opensource repository ElegantRL [^3] https://github.com/AI4Finance-Foundation/ElegantRL. The explore rate of the DQN network is 0.5, and the discount factor of the future rewards is 0.98. The middle layer dimension of MLP is set to [80, 40]. The weight for both semantic similarity and perceptual similarity is set to 0.5, i.e., $a=b=0.5$. $\lambda$ and $\delta$ in the reward function of the inner budget allocation is set to 0.5.
Implementation Settings. We have 8 data holders with different amount (50, 60, 80, 100, 150, 200) and different quality (low, medium, high) of training data in the system. Each data holder will have a self-proposed price for their data, which is publicly known. When the allocated payment meet the self-proposed price, the data holder will join the training for the current round.
Evaluation Metrics. We use a modified version of Fréchet Inception Distance (FID) score (Heusel et al., 2018) $Q(\mathcal{M})$ to measure the output quality of diffusion models:
(23) $Q(\mathcal{M}):=\frac{100}{FID+10^{-6}}.$
FID measures how similar the distribution of generated images is to the distribution of real images, based on features extracted by a pre-trained Inception network. The formulation of the metric is as follows:
(24) $\text{FID}=\|m_{x}-m_{y}\|^{2}+\text{Tr}(\text{Cov}(x)+\text{Cov}(y)-2(\text{Cov}(x)\text{Cov}(y))^{1/2})$
where: $m_{x}$ and $\text{Cov}(x)$ represent the mean and covariance matrix of the feature vectors from real images; $m_{y}$ and $\text{Cov}(y)$ represent the mean and covariance matrix of the feature vectors from generated images; $\text{Tr}(\cdot)$ denotes the trace of a matrix.
(a) ArtBench
(b) Portrait
(c) Cartoon
Figure 5. Ablation Results on Three Datasets.
Baseline. We integrate three strategies: random, linear, and greedy strategies to the experiments and compare the proposed method with 10 different baselines:
  1. (1)
    Greedy + Linear(G+L) uses greedy method for outer layer budget distribution, and the linear method for inner layer budget distribution;
  2. (2)
    Linear + Linear(L+L) linearly distributes the budget at the outer and inner layer;
  3. (3)
    Random + Random(R+R) randomly distributes the budget at the outer and inner layer;
  4. (4)
    RL + Random(RL+R) uses RL for outer layer budget distribution, and randomly distributes the budget at the inner layer;
  5. (5)
    RL + Linear(RL+L) uses RL for outer layer budget distribution, and linearly distributes the budget at the inner layer;
  6. (6)
    Random + RL(R+RL) randomly distributes the budget at the outer layer, and uses the RL for inner distribution;
  7. (7)
    Linear + RL(L+RL) linearly distributes the budget at the outer layer, and uses RL for inner distribution;
  8. (8)
    Greedy + RL(G+RL) uses greedy method for outer layer budget distribution, and uses the RL for inner distribution.
5.2. Results and Discussion
The comparision results are shown in Table 2 and Figure. 4. It can be observed that on all the three datasets, our method consistently outperform the baselines in terms of the model quality. Our proposed method can optimize the budget distribution and maximize the quality of the model, which validates the effectiveness of the proposed method.
We can also find that
  1. (1)
    Methods using RL generally outperforms other methods in terms of the model quality. Most top-performing approaches integrate RL, as RL has strong abilities of finding optimal budget distributions. Our method with hierarchical RL structure can optimize the outcome from a global perspective and achieve the highest performance.
  2. (2)
    Among the methods using single layer RL, methods using RL at the outer layer outperform the methods using RL at the inner layer in terms of model quality. The results imply that the global decisions that the outer RL deals with might have a larger impact on the overall model performance.
  3. (3)
    Methods using linear distribution in both layers have bad performance in terms of model quality, since uniform allocation does not capture differences between data holders and between rounds.
  4. (4)
    Methods combining random strategy are generally not well-performed. The random strategy does no optimization for the system, and it has a considerably large variance comparing to the other methods due to the randomness.
5.2.1. Ablation Study
We conduct ablation experiments to explore the impact of only considering the contributions of data participants and copyright loss on the results.
  1. (1)
    RL + Contribution(RL+C) uses RL for outer distribution, and distributes the budget according to the contribution of the data holder at the inner layer;
  2. (2)
    RL + Copyright loss(RL+CL) uses RL for outer distribution, and distributes the budget according to the copyright loss of the data holder at the inner layer;
As shown in Table 3, our method outperforms both the ablated versions in terms of model quality. The results show that the contribution and copyright loss are both playing crucial roles in enhancing the model performance on efficiently allocate the budget.
Table 3. Ablation Studies on Three Datasets
Dataset Ours RL+C RL+CL
ArtBench 6.178$\pm$0.344 5.102$\pm$0.251 5.284$\pm$0.121
Portrait 8.331$\pm$1.785 6.759$\pm$0.764 6.390$\pm$0.695
Cartoon 6.038$\pm$0.737 5.624$\pm$0.490 5.741$\pm$0.616
6. Conclusion and Future Work
This paper introduces a novel method to address the challenge of copyright infringement in generative art with an economic method. Our approach introduces a novel copyright metric, grounded in legal perspective, to evaluate copyright loss and integrates the TRAK method to quantify the contribution of data holders. Through a hierarchical budget allocation method using reinforcement learning, we demonstrate how remuneration can be fairly distributed based on each data holder’s contribution and copyright impact. Extensive experiments validate the efficacy of this scheme, significantly outperforming existing incentive scheme in terms of model quality.
In the future, we plan to expedite the calculation of copyright loss and contributions to adapt the model for dynamic environments where data holders may join and leave. Additionally, current methods that assign identical copyright losses to duplicated samples can result in repeated payments for the same content. This issue calls for further exploration to refine the handling of duplicates in datasets.